# DiFlowDubber: Discrete Flow Matching for Automated Video Dubbing via Cross-Modal Alignment and Synchronization

### 17 Mar 2026 v2

@@ 4.3 Qualitative Analysis
Figure 3: Mel-spectrogram visualization compared with Ground Truth (GT) speech.
Figure 3 provides a qualitative comparison of the generated mel-spectrograms across different dubbing systems. Our method exhibits a strong resemblance to the ground-truth recording. The highlighted regions (in red boxes) indicate areas where different models show noticeable discrepancies, with our results remaining most consistent with the ground truth. This demonstrates that our approach effectively preserves pitch continuity and prosodic dynamics while maintaining alignment with lip movements in the corresponding video frames. Overall, these results confirm that our model produces speech with more natural rhythm, precise synchronization, and expressive delivery that closely mirrors human dubbing performance.
@@ 4.4 Ablation Study
As shown in Table 3, we observe that removing the $\mathcal{L}_{\text{distill}}$ loss (see row #6) leads to degraded pronunciation accuracy, demonstrating the importance of effectively transferring knowledge from the TTS domain, where pronunciation is highly accurate, to the video dubbing task while maintaining temporal lip-sync alignment. When the $\mathcal{L}_{\text{CTC}}$ loss is removed (see row #7), a similar pattern is observed, highlighting its role in facilitating more accurate alignment learning between the input text and the output of the Synchronizer. From row #8, fusing global prosody prior with content information enhances speaker similarity, indicating that face-derived global prosody encodes speaker-dependent cues. While these variants improve lip synchronization, they also introduce trade-offs in pronunciation accuracy, speaker similarity, and naturalness.
@@
Table 3 shows that removing the FaPro module (see row #9) results in consistent quality degradation across almost all metrics, confirming the importance of global prosody cues derived from facial expressions. These results indicate that FaPro effectively captures expressive facial dynamics to guide prosody modeling.
@@ 5 Conclusion
In this paper, we propose DiFlowDubber, a novel end-to-end video dubbing system that is the first to extend discrete flow matching to the field of video dubbing. Our training framework consists of two stages. First, we pre-train a zero-shot TTS model on a large-scale TTS corpus, resulting in a model that can capture linguistic information, expressive prosody and realistic acoustic characteristics. Subsequently, video adaptation is performed on the TTS model. In this stage, we introduce the Synchronizer module with a dual-alignment mechanism, achieving precise synchronization between all three modalities video, text, and speech. Both quantitative and perceptual evaluations confirm the superior performance of our method compared with prior approaches.
@@ B.2.1 Zero-shot TTS Pre-training Stage
Denoiser Architecture
We use a Diffusion Transformer (DiT) [43] as the denoiser in our discrete flow matching framework. The DiT consists of 8 layers with a hidden size of 768 and 8 attention heads, and is enhanced with rotary position embeddings (RoPE) [54]. The time step is embedded using an MLP that maps it to a 768-dimensional vector, while the speaker embedding, originally of dimension 256, is linearly projected to the same 768-dimensional space. The global conditioning vector is then formed by summing the resulting time and speaker embeddings, which are used to condition the adaptive layer normalization (AdaLN) layers. We employ a long skip connection from the input to the output of the final Transformer block. The final prediction stage performs a non-affine layer normalization followed by AdaLN modulation and a linear projection. The global conditioning vector is processed through a SiLU-activated [16] MLP to generate shift and scale parameters that modulate the normalized hidden states. The final linear projection produces an output of $(1+3)\times 768$ dimensions, corresponding to one prosody quantizer and three acoustic quantizers.
Contextual Modeling
We describe the construction of the input to the Denoiser. The input consists of two parts: the conditioning input, denoted as $\mathbf{e}_{\text{cond}}$, and the current denoising target, $\mathbf{e}_{t}$.
For $\mathbf{e}_{\text{cond}}$, given a reference speech sample, we use FACodec [28] to extract discrete speech tokens along with a speaker embedding, yielding $\tilde{\mathbf{x}}^{p}\in\mathbb{R}^{m\times L^{\prime}}$, $\tilde{\mathbf{x}}^{c}\in\mathbb{R}^{n\times L^{\prime}}$, $\tilde{\mathbf{x}}^{a}\in\mathbb{R}^{k\times L^{\prime}}$, and $\mathbf{s}\in\mathbb{R}^{D_{s}}$, representing the prosody, content, and acoustic tokens, and the speaker embedding, respectively. Here, $m$, $n$, and $k$ denote the number of residual vector quantization (RVQ) codebooks for prosody, content, and acoustic representations, each with a vocabulary size of $v$. $L^{\prime}$ is the token sequence length, and $D_{s}$ is the speaker embedding dimension. We pass $\tilde{\mathbf{x}}_{p}$ and $\tilde{\mathbf{x}}_{a}$ to dedicated embedders to encode them into latent representations $\tilde{\mathbf{e}}_{p}\in\mathbb{R}^{m\times L^{\prime}\times D}$ and $\tilde{\mathbf{e}}_{a}\in\mathbb{R}^{k\times L^{\prime}\times D}$. In the zero-shot TTS setting, which aims to mimic the speaker style of the reference speech, we omit the content tokens $\tilde{\mathbf{x}}_{c}$ and replace them with a zero embedding $\tilde{\mathbf{e}}_{\text{zeros}}\in\mathbb{R}^{n\times L^{\prime}\times D}$. The final conditioning vector is defined as $\mathbf{e}_{\text{cond}}=\tilde{\mathbf{e}}_{p}\oplus\tilde{\mathbf{e}}_{\text{zeros}}\oplus\tilde{\mathbf{e}}_{a}$, where $\mathbf{e}_{\text{cond}}\in\mathbb{R}^{(m+n+k)\times L^{\prime}\times D}$, and $\oplus$ denotes the concatenation operation.
For $\mathbf{e}_{t}$, given the current denoising target tokens $\mathbf{x}_{t}=\mathbf{x}_{t}^{p}\oplus\mathbf{x}_{t}^{a}$, we feed $\mathbf{x}_{t}^{p}\in\mathbb{R}^{m\times L}$ and $\mathbf{x}_{t}^{a}\in\mathbb{R}^{k\times L}$ into the same dedicated embedders to obtain latent representations $\mathbf{e}_{t}^{p}\in\mathbb{R}^{m\times L\times D}$ and $\mathbf{e}_{t}^{a}\in\mathbb{R}^{k\times L\times D}$. Since speech rhythm also depends on linguistic content [66], we incorporate the content information from the final hidden state of the content modeling module, denoted as $\mathbf{e}_{c}$. The final denoising target is then defined as $\mathbf{e}_{t}=\mathbf{e}_{t}^{p}\oplus\mathbf{e}_{c}\oplus\mathbf{e}_{t}^{a}$, where $\mathbf{e}_{t}\in\mathbb{R}^{(m+n+k)\times L\times D}$.
Finally, the complete input to the Denoiser is constructed by concatenating $\mathbf{e}_{\text{cond}}$ and $\mathbf{e}_{t}$ along the temporal dimension: $\mathbf{e}=\mathbf{e}_{\text{cond}}\oplus\mathbf{e}_{t}$, where $\mathbf{e}\in\mathbb{R}^{(m+n+k)\times(L^{\prime}+L)\times D}$. The resulting tensor $\mathbf{e}$ is then permuted to shape $\mathbb{R}^{(L^{\prime}+L)\times(m+n+k)D}$ and projected to $\mathbb{R}^{(L^{\prime}+L)\times D}$ before being fed into the Denoiser.
@@
The objective of Discrete Flow Matching (DFM) is to transform source samples $\mathbf{x}_{0}\sim p$ into target samples $\mathbf{x}_{1}\sim q$. In our formulation, the source distribution corresponds to all-mask tokens [MASK], while the target distribution $\mathbf{x}_{1}=\mathbf{x}_{1}^{p}\oplus\mathbf{x}_{1}^{a}$ is factorized into prosodic $\mathbf{x}_{1}^{p}$ and acoustic $\mathbf{x}_{1}^{a}$ components, enabling structured joint learning. Following [19], we employ a monotonic scheduler $\kappa_{t}\in[0,1]$ with boundary conditions $\kappa_{0}=0$ and $\kappa_{1}=1$, where $t\in[0,1]$ denotes continuous time. In our implementation, we set $\kappa_{t}=t^{2}$. This scheduler progressively interpolates between the source and target distributions, smoothly transitioning from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ as $\kappa_{t}$ increases. We then construct a conditional probability path, referred to as the mixture path, which linearly interpolates between the source and target distributions: $p_{t}(\mathbf{x}^{i}|\mathbf{x}_{0},\mathbf{x}_{1})=(1-\kappa_{t})\,p_{0}(\mathbf{x}^{i}|\mathbf{x}_{0})+\kappa_{t}\,p_{1}(\mathbf{x}^{i}|\mathbf{x}_{1}).$ This path defines an evolution of discrete states governed by a probability velocity field $\mathbf{u}_{t}$, expressed as:
@@
The overall training objective in this stage consists of three components. The first term, $\mathcal{L}_{d}$, is optimized using mean squared error (MSE). The second term, $\mathcal{L}_{c}$, corresponds to the content modeling loss, computed as the cross-entropy between the predicted logits and the target content sequence. The third term is the discrete flow-matching loss, $\mathcal{L}_{\text{DFM}}$. The total loss is formulated as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{d}+\lambda_{2}\mathcal{L}_{c}+\lambda_{3}\mathcal{L}_{\text{DFM}},$ (9)
@@ B.2.2 Video Dubbing Adaptation Stage
The architecture is composed of a learnable upsampling layer, a ConvNeXt-V2 [59] encoder stack, layer normalization, and a Transformer decoder. The upsampling layer applies linear interpolation followed by 4 sets of learnable convolutional transformations. Each set consists of a 1D convolution (kernel size 3, stride 1, padding 1), a group normalization layer with 1 group, and a Mish activation [40] function. A final 1D convolution projects the features to the target hidden dimension of 256. The upsampled features are then processed by a stack of 8 ConvNeXt-V2 [59] blocks, each with hidden dimension 256 and intermediate dimension 1024. Every block contains a depthwise 1D convolution with kernel size 7, followed by layer normalization and 2 linear projections. Between these projections, we apply a GELU activation [23] and a Global Response Normalization (GRN) module, which stabilizes the feature scale by normalizing the response of each channel with respect to its global L2 magnitude. A residual connection is added around each block, enabling stable and efficient learning of temporal dynamics across long facial sequences. After the ConvNeXt-V2 [59] encoder stack, a layer normalization layer is applied before forwarding the features to one Transformer decoder that serves as the Prosody Predictor, built on the FFT block architecture. The decoder uses 4 attention heads, a hidden dimension of 256, and a feed-forward expansion ratio of 4, and operates with a maximum sequence length of 5000.
@@
As shown in Figure 4, the module consists of 2 Transformer stacks, a learnable upsampling layer, a light fusion network, a ConvNeXt-V2 [59] encoder stack, and two layer normalization layers. We use 8 Transformer blocks for both video-text and speech-text alignment, where the attention mechanism follows a Monotonic Multi-Head Attention [36] formulation with 4 attention heads, a hidden dimension of 256, and an intermediate dimension of 1024. The learnable upsampling layer and the ConvNeXt-V2 [59] encoder stack use the same architecture and hyperparameters as those in the FaPro module. The light fusion network is implemented as a linear projection layer.
@@ B.3 Training Details
Training is conducted in 2 stages:
@@
We train on 470 hours of the LibriTTS dataset [62] using 4 NVIDIA A100 GPUs with a batch size of 16 for a total of 300k training steps. The learning rate is set to $10^{-4}$ with a 10% warm-up ratio and a weight decay of 0.01, optimized using the AdamW optimizer [35]. The weighting coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are defined as in Equation 9.
@@
Training is performed on the Chem [25] and GRID [15] datasets using 4 NVIDIA A100 GPUs with batch sizes of 16 and 64, for 150 and 200 epochs, respectively. We use the AdamW optimizer [35] with a learning rate of $10^{-3}$, a weight decay of 0.01, and a 5% warm-up ratio. The weights for the total loss $\mathcal{L}_{\text{total}}$ (defined in the main paper) are set to $\lambda_{1}=1.0$, $\lambda_{2}=0.1$, $\lambda_{3}=0.1$, $\lambda_{4}=1.0$, $\lambda_{5}=0.001$, and $\lambda_{6}=0.001$.

### 03 Apr 2026 v4

@@ 4.3 Qualitative Analysis
Figure 3: Mel-spectrogram visualization compared with Ground-Truth speech.
Figure 3 provides a qualitative comparison of the mel-spectrograms generated in different dubbing systems. The highlighted regions (in red boxes) indicate areas where different models show noticeable discrepancies in speech quality. Across all samples, DiFlowDubber produces spectrograms most similar to the ground truth, preserving clear harmonic structure, accurate temporal boundaries, and natural prosodic dynamics. In addition, it effectively maintains pitch continuity and alignment with lip movements in the corresponding video frames, confirming that our model produces speech with more natural rhythm, precise synchronization, and expressive delivery.
@@ 4.4 Ablation Study
As shown in Table 3, we observe that removing the $\mathcal{L}_{\text{distill}}$ loss (see row #6) leads to degraded pronunciation accuracy, demonstrating the importance of effectively transferring knowledge from the TTS domain, where pronunciation is highly accurate, to the video dubbing task while maintaining temporal lip-sync alignment. When the $\mathcal{L}_{\text{CTC}}$ loss is removed (see row #7), a similar pattern is observed, highlighting its role in facilitating more accurate alignment learning between the input text and the output of the Synchronizer. From row #8, fusing global prosody prior with content information enhances speaker similarity, indicating that face-derived global prosody encodes speaker-dependent cues essential for perceptually consistent dubbing.
@@
Table 3 shows that removing the FaPro module (see row #9) results in consistent quality degradation across almost all metrics, confirming the importance of global prosody cues derived from facial expressions. These results indicate that FaPro effectively captures expressive facial dynamics to guide prosody generation.
@@ 5 Conclusion
In this paper, we present DiFlowDubber, the first end-to-end video dubbing framework built upon a discrete flow matching backbone. Our two-stage pipeline first pre-trains a zero-shot TTS system, where a simple deterministic architecture captures linguistic structure and the DFPA module models expressive prosody and realistic acoustic characteristics from large-scale TTS corpora. In the video dubbing adaptation stage, the CCTA module transfers content knowledge from the TTS domain via the Synchronizer, which enforces precise cross-modal synchronization through a dual-alignment mechanism over text, video, and speech. Meanwhile, the FaPro extracts global prosodic cues from facial expressions. These cues, together with the outputs of the Synchronizer, are fused to capture prosody-content correlations, thereby guiding the DFPA to generate expressive prosody and acoustic tokens for content-consistent speech. Extensive experimental results validate the effectiveness of our approach.
@@ B.2.1 Zero-shot TTS Pre-training Stage
The objective of Discrete Flow Matching is to transform source samples $\mathbf{x}_{0}\sim p$ into target samples $\mathbf{x}_{1}\sim q$. In our formulation, the source distribution corresponds to all-mask tokens [MASK], while the target distribution $\mathbf{x}_{1}=[\mathbf{x}_{p,1};\mathbf{x}_{a,1}]\in\mathbb{N}^{(m+k)\times L}$ is factorized into prosodic $\mathbf{x}_{p,1}\in\mathbb{N}^{m\times L}$ and acoustic $\mathbf{x}_{a,1}\in\mathbb{N}^{k\times L}$ components, enabling structured joint learning. Following [18], we employ a monotonic scheduler $\kappa_{t}\in[0,1]$ with boundary conditions $\kappa_{0}=0$ and $\kappa_{1}=1$, where $t\in[0,1]$ denotes continuous time. In our implementation, we set $\kappa_{t}=t^{2}$. This scheduler progressively interpolates between the source and target distributions, smoothly transitioning from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ as $\kappa_{t}$ increases. We then construct a conditional probability path, referred to as the mixture path, which linearly interpolates between the source and target distributions: $p_{t}(\mathbf{x}^{i}|\mathbf{x}_{0},\mathbf{x}_{1})=(1-\kappa_{t})\,p_{0}(\mathbf{x}^{i}|\mathbf{x}_{0})+\kappa_{t}\,p_{1}(\mathbf{x}^{i}|\mathbf{x}_{1}).$ This path defines an evolution of discrete states governed by a probability velocity field $\mathbf{u}_{t}$, expressed as:
@@
Denoiser Architecture
We use a Diffusion Transformer (DiT) [41] as the denoiser in our discrete flow matching framework. The DiT consists of 8 layers with a hidden size of 768 and 8 attention heads, and is enhanced with rotary position embeddings (RoPE) [51]. The time step is embedded using an MLP that maps it to a 768-dimensional vector, while the speaker embedding, originally of dimension 256, is linearly projected to the same 768-dimensional space. The global conditioning vector $\mathbf{c}_{g}$ is then formed by summing the resulting the time embedding $\mathbf{t}$ and the speaker embedding $\mathbf{s}$ projected via an MLP, which are used to condition the Adaptive Layer Normalization (AdaLN) layers. We employ a long skip connection from the input to the output of the final Transformer block. The final prediction stage performs a non-affine layer normalization followed by AdaLN modulation and a linear projection. The global conditioning vector is processed through a SiLU-activated [15] MLP to generate shift and scale parameters that modulate the normalized hidden states. The final linear projection produces an output of $(1+3)\times 768$ dimensions, corresponding to one prosody quantizer and three acoustic quantizers.
Contextual Modeling
We describe the construction of the input to the Denoiser. The input consists of two parts: the conditioning input, denoted as $\mathbf{e}_{\text{cond}}$, and the current denoising target, $\mathbf{e}_{t}$ at time step $t$.
For $\mathbf{e}_{\text{cond}}$, given a reference speech sample $\mathbf{S}^{r}$, we use FACodec [27] to extract discrete speech tokens along with a speaker embedding, yielding $\mathbf{x}^{r}_{p}\in\mathbb{N}^{m\times L^{\prime}}$, $\mathbf{x}^{r}_{c}\in\mathbb{N}^{n\times L^{\prime}}$, $\mathbf{x}^{r}_{a}\in\mathbb{N}^{k\times L^{\prime}}$, and $\mathbf{s}\in\mathbb{R}^{D_{s}}$, representing the prosody, content, and acoustic token sequences, and the speaker embedding, respectively. Here, $m$, $n$, and $k$ denote the number of Residual Vector Quantization (RVQ) codebooks for prosody, content, and acoustic representations, each with a vocabulary size of $v$. $L^{\prime}$ is the token sequence length, and $D_{s}$ is the speaker embedding dimension. We pass $\mathbf{x}^{r}_{p}$ and $\mathbf{x}^{r}_{a}$ to dedicated embedders to encode them into latent representations $\mathbf{e}^{r}_{p}\in\mathbb{R}^{m\times L^{\prime}\times D}$ and $\mathbf{e}^{r}_{a}\in\mathbb{R}^{k\times L^{\prime}\times D}$. In the zero-shot TTS setting, which aims to mimic the speaker style of the reference speech, we omit the content tokens $\mathbf{x}^{r}_{c}$ and replace them with zero embeddings. The final conditioning vector is defined as $\mathbf{e}_{\text{cond}}=[\mathbf{e}^{r}_{p};\mathbf{0};\mathbf{e}^{r}_{a}]\in\mathbb{R}^{(m+n+k)\times L^{\prime}\times D}$.
For $\mathbf{e}_{t}$, given the current denoising target tokens $\mathbf{x}_{t}=[\mathbf{x}_{p,t};\mathbf{x}_{a,t}]\in\mathbb{N}^{(m+k)\times L}$, we feed $\mathbf{x}_{p,t}\in\mathbb{N}^{m\times L}$ and $\mathbf{x}_{a,t}\in\mathbb{N}^{k\times L}$ into the dedicated embedders to obtain latent representations: $\mathbf{e}_{p,t}=\mathbf{E}_{p}(\mathbf{x}_{p,t})\in\mathbb{R}^{m\times L\times D}$ and $\mathbf{e}_{a,t}=\mathbf{E}_{a}(\mathbf{x}_{a,t})\in\mathbb{R}^{k\times L\times D}$. Since speech rhythm also depends on linguistic content [63], we incorporate the content information from the final hidden state of the content modeling module, denoted as $\tilde{\mathbf{h}}_{c}$. The final denoising target is then defined as $\mathbf{e}_{t}=[\mathbf{e}_{p,t};\tilde{\mathbf{h}}_{c};\mathbf{e}_{a,t}]\in\mathbb{R}^{(m+n+k)\times L\times D}$.
Finally, the complete input to the Denoiser is constructed by concatenating $\mathbf{e}_{\text{cond}}$ and $\mathbf{e}_{t}$ along the temporal dimension: $\mathbf{e}=[\mathbf{e}_{\text{cond}};\mathbf{e}_{t}]\in\mathbb{R}^{(m+n+k)\times(L^{\prime}+L)\times D}$. The input $\mathbf{e}$ is permuted to shape $\mathbb{R}^{(L^{\prime}+L)\times(m+n+k)D}$, projected to $\mathbb{R}^{(L^{\prime}+L)\times D}$, and fed into the Denoiser. The resulting output is then passed through a final transformation layer comprising layer normalization, AdaLN-based modulation conditioned on $\mathbf{c}_{g}$, and a linear projection, yielding features of shape $\mathbb{R}^{(L^{\prime}+L)\times(m+k)D}$. We discard the conditioning portion and permute the result to obtain the final hidden representation in $\mathbb{R}^{(m+k)\times L\times D}$, which is then passed through a dedicated MLP to produce the posterior distributions over prosody and acoustic tokens in $\mathbb{R}^{(m+k)\times L\times v}$.
@@
The overall training objective in this stage consists of three components. First, we optimize the Duration Predictor using the Mean Squared Error loss on the logarithmic scale, denoted as $\mathcal{L}_{d}$. Second, $\mathcal{L}_{c}$ corresponds to the content modeling loss, computed as the Cross-Entropy between the predicted logits and the target content sequence. Third, we employ the discrete flow matching loss, $\mathcal{L}_{\text{DFM}}$ (described in the main paper). The overall loss is formulated as follow:
$\mathcal{L}_{\text{TTS}}=\lambda_{1}\mathcal{L}_{d}+\lambda_{2}\mathcal{L}_{c}+\lambda_{3}\mathcal{L}_{\text{DFM}},$ (9)
@@ B.2.2 Video Dubbing Adaptation Stage
The architecture is composed of a learnable upsampling layer, a ConvNeXt V2 [56] encoder stack, layer normalization, and a Transformer decoder. The upsampling layer applies linear interpolation followed by 4 sets of learnable convolutional transformations. Each set consists of a 1D convolution (kernel size 3, stride 1, padding 1), a group normalization layer with 1 group, and a Mish activation [38] function. A final 1D convolution projects the features to the target hidden dimension of 256. The upsampled features are then processed by a stack of 8 ConvNeXt V2 [56] blocks, each with hidden dimension 256 and intermediate dimension 1024. Every block contains a depthwise 1D convolution with kernel size 7, followed by layer normalization and 2 linear projections. Between these projections, we apply a GELU activation [22] and a Global Response Normalization (GRN) module, which stabilizes the feature scale by normalizing the response of each channel with respect to its global L2 magnitude. A residual connection is added around each block, enabling stable and efficient learning of temporal dynamics across long facial sequences. After the ConvNeXt V2 [56] encoder stack, a layer normalization layer is applied before forwarding the features to one Transformer decoder that serves as the Prosody Predictor, built on the FFT block architecture. The decoder uses 4 attention heads, a hidden dimension of 256, and a feed-forward expansion ratio of 4, and operates with a maximum sequence length of 5000.
@@
As shown in Figure 4, the module consists of 2 Transformer stacks, a learnable upsampling layer, a light fusion network, a ConvNeXt V2 [56] encoder stack, and two layer normalization layers. We use 8 Transformer blocks for both video-text and speech-text alignment, where the attention mechanism follows a Monotonic Multi-Head Attention [34] formulation with 4 attention heads, a hidden dimension of 256, and an intermediate dimension of 1024. The learnable upsampling layer and the ConvNeXt V2 [56] encoder stack use the same architecture and hyperparameters as those in the FaPro module. The light fusion network is implemented as a linear projection layer.
@@ B.3 Training Details
The training is conducted in 2 stages:
@@
We train on 470 hours of the LibriTTS dataset [59] using 4 NVIDIA A100 80GB GPUs with a batch size of 16 for a total of 300k training steps. The learning rate is set to $10^{-4}$ with a 10% warm-up ratio and a weight decay of 0.01, optimized using the AdamW optimizer [33]. The weighting coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are defined as in Equation 9.
@@
Training is performed on the Chem [24] and GRID [14] datasets using 4 NVIDIA A100 80GB GPUs with batch sizes of 16 and 64, for 150 and 200 epochs, respectively. We use the AdamW optimizer [33] with a learning rate of $10^{-3}$, a weight decay of 0.01, and a 5% warm-up ratio. The weights for the loss $\mathcal{L}_{\text{Dubbing}}$ (defined in the main paper) are set to $\lambda_{1}=1.0$, $\lambda_{2}=0.1$, $\lambda_{3}=0.1$, $\lambda_{4}=1.0$, $\lambda_{5}=0.001$, and $\lambda_{6}=0.001$.